# From Polysemy to Quantum Attention: Context, Randomness, and Hilbert Space Perspectives on Large Language Models

### v3

@@ Abstract
We trace a conceptual path from linguistic ambiguity to quantum-inspired architectures in the study of Large Language Models (LLMs). First, we examine polysemy and show how modern transformer networks resolve multiple word meanings through dynamic, context-dependent embeddings, contrasting this with an alternative framework rooted in quantum contextuality. Next, we consider stochasticity not as a flaw but as a central design principle: randomness in sampling provides the mechanism by which LLMs generate diverse, creative outputs and explore their own representational landscape. Finally, we introduce the notion of quantum attention, a speculative extension of the classical attention mechanism that constrains its transformations to be unitary. This shift reframes attention as a reversible rotation in Hilbert space rather than a dissipative mixture, aligning the processing of meaning in language models with fundamental structures of quantum theory.
@@ I Introduction
This work moves beyond a procedural overview to focus on three foundational aspects. We will address the model’s ability to handle polysemy in Section VI, exploring how modern architectures dynamically construct context-aware vector representations in a departure from older, static methods.
In Section VII, we will investigate the origin of stochasticity in their responses, framing the probabilistic nature of text generation as an intentional and crucial feature that underpins their creativity and flexibility.
Finally, in Section IX we will draw a compelling, albeit analogical, comparison to the formalism of quantum mechanics, using the shared mathematical language of vector spaces and context-dependent projections to offer a novel framework for conceptualizing how these systems process information and construct meaning. By dissecting these three pillars, this note aims to provide a more nuanced perspective on the inner workings of Large Language Models.
@@ V Reinforcement Learning: Refining Reasoning
VI Polysemy
@@
VI.1 Dynamic Contextualization via Self-Attention
@@
VI.2 An Alternative Framework: Quantum Contextuality
@@
VII The Origin of Stochasticity in responses
@@
VIII Algorithmic Capacity and the Role of Stochastic Seeding
@@ VIII.1 The LLM as a Formal System
VIII.2 Stochasticity as an Engine for Discovery
@@
VIII.3 Comparison to Genetic Algorithms
@@
IX Comparison to Quantum Mechanics
@@
IX.1 Vectors, Inner Products, and Contexts
@@
IX.2 Dynamics: Unitary Evolution vs. Transformer Block Updates
@@
IX.3 Contextuality in Both Senses
@@
IX.4 Operationalizing an Internal Observer
@@
IX.5 Orthogonal Bases and Functional Decomposition
@@
IX.6 Analogies, Disanalogies, and Conclusion
@@ X Quantum Attention
X.1 Goals of Quantum Attention
@@
X.2 Constraints and Learning Mechanisms
@@
X.3 Comparison with Classical Attention
@@ Appendix C The Self-Attention Mechanism
Let us consider a sequence of $n$ input tokens. Each token is represented by a learned embedding vector, and these row vectors are stacked to form an input matrix $X\in\mathbb{R}^{n\times d}$, where $d$ is the dimension of the embedding space (often called $d_{\text{model}}$ in literature). The self-attention mechanism computes an output matrix $Y\in\mathbb{R}^{n\times d}$ where each output row $\bm{y}_{i}$ is a new, contextualized representation of the corresponding input token $\bm{x}_{i}$. This is achieved through the following steps.
@@ C.1 Projection to Query, Key, and Value
First, the input matrix $X$ is linearly projected into three distinct matrices: Query ($\bm{Q}$), Key ($\bm{K}$), and Value ($\bm{V}$). This is accomplished using three learned weight matrices: $\bm{W}_{Q}\in\mathbb{R}^{d\times d_{k}}$, $\bm{W}_{K}\in\mathbb{R}^{d\times d_{k}}$, and $\bm{W}_{V}\in\mathbb{R}^{d\times d_{v}}$. Typically, the dimensions are set such that $d_{k}=d_{v}$.
@@ D.1 Multi-Head Attention
    Projection: Instead of one set of projection matrices, the model learns $h$ independent sets of matrices ($\bm{W}_{Q,i},\bm{W}_{K,i},\bm{W}_{V,i}$ for each head $i=1,\dots,h$). The input $X$ is projected through each of these sets, creating $h$ different Query, Key, and Value representations.
@@
    Concatenation and Final Projection: The outputs of these $h$ heads are concatenated along the feature dimension into a single, larger matrix. This matrix is then projected back to the original model dimension $d$ using an additional learned weight matrix, $\bm{W}_{O}\in\mathbb{R}^{hd_{v}\times d}$, which learns to aggregate the information from all heads.
    $\text{MultiHead}(X)=\text{Concat}(\text{head}_{1},\dots,\text{head}_{h})\bm{W}_{O}.$ (45)

### v4

# Perspectives on Large Language Models: Polysemy, Stochasticity, Exponential Expressibility, and Unitary Attention

@@ Abstract
This paper explores foundational aspects of Large Language Models (LLMs). We analyze how the expressibility of semantic features scales exponentially with embedding space dimensions using quasi-orthogonal vectors. We contrast the dynamic, context-dependent embeddings of Transformer architectures, which resolve polysemy, with a static vector approach based on quantum contextuality. Stochasticity is framed as an essential feature for enabling creative output through probabilistic sampling. Finally, we propose quantum attention as a unitary extension of classical mechanisms, reframing LLM processing as reversible, quantum-like evolutions in Hilbert space.
@@ I Introduction
This work moves beyond a procedural overview to focus on three foundational aspects. We will address the model’s ability to handle polysemy in Section VIII, exploring how modern architectures dynamically construct context-aware vector representations in a departure from older, static methods.
In Section IX, we will investigate the origin of stochasticity in their responses, framing the probabilistic nature of text generation as an intentional and crucial feature that underpins their creativity and flexibility.
Finally, in Section XI we will draw a compelling, albeit analogical, comparison to the formalism of quantum mechanics, using the shared mathematical language of vector spaces and context-dependent projections to offer a novel framework for conceptualizing how these systems process information and construct meaning. By dissecting these three pillars, this note aims to provide a more nuanced perspective on the inner workings of Large Language Models.
@@ V Reinforcement Learning: Refining Reasoning
VI Self-Attention and Computational Universality
The apparent self-reflection observed in Large Language Models (LLMs), often manifested through phrases such as “let me reconsider” or “wait, the user implies…”, does not emerge from conscious introspection but is rather a learned artifact of the training data. These models are trained to predict the next token in a sequence, and such reflective phrases are statistically probable continuations in contexts that necessitate re-evaluation, thereby mimicking human cognitive patterns.
The core mechanism enabling this behavior is the self-attention layer within the Transformer architecture. Self-attention allows the model to weigh the importance of all tokens in the preceding sequence—including its own generated output—at each computational step. This process is enhanced by multi-head attention, which permits the model to track disparate linguistic and logical relationships simultaneously in parallel subspaces. Therefore, while self-attention provides the mechanical capability for in-context reflection, this behavior is fundamentally a stylistic pattern. More profoundly, this same mechanism is the foundation for the model’s capacity for universal computation.
This computational capacity of LLMs extends beyond mere linguistic processing, touching upon the theoretical limits of computability. Transformer-based models possess the intrinsic capability to compute partial recursive functions, thereby positioning them as computationally universal systems analogous to a Universal Turing Machine (UTM). This equivalence is most clearly conceptualized under the assumption of an infinitely expandable context window, which would function as the unbounded tape of a UTM. However, the practical technological constraints of current hardware and model architectures mean that this computational universality remains more of a theoretical attribute than a fully realized, practical reality.
The computational universality of LLMs is not explicitly programmed but emerges from the architecture’s design. Working within an autoregressive decoding loop, the self-attention mechanism allows the model to maintain and update its state by attending to its own previous outputs. This process is powerful enough to simulate known Turing-complete systems, such as a Lag system [6], effectively using the generated text as a computational tape.
The primary limitation to this capability is the model’s finite context window, a direct consequence of the quadratic computational and memory scaling of the self-attention mechanism with respect to sequence length. Consequently, like any physical computing device, an LLM is constrained by finite memory, precluding true mathematical Turing completeness, which presupposes infinite storage.
In practice, LLMs demonstrate a nuanced aptitude for handling recursion, particularly within the domain of code generation. This holistic reasoning is facilitated by self-attention, which permits the model to cross-reference different parts of the generated code, ensuring that recursive calls correctly relate to their corresponding base cases and termination conditions. Such capabilities serve as practical evidence of the sophisticated, Turing-complete computations that self-attention enables.
VII Exponential Scaling of Expressive Capacity via Quasi-Dimensionality
The dense vector representations, or embeddings, of tokens in Large Language Models (LLMs) are foundational to their ability to capture complex semantic relationships. These relationships can often be revealed through simple vector arithmetic. A canonical example involves the embeddings for woman and man, denoted as $\bm{v}(\text{woman})$ and $\bm{v}(\text{man})$, respectively. The difference between these vectors, $\bm{v}(\text{woman})-\bm{v}(\text{man})$, yields a new vector that points in a semantic direction associated with gender. This direction is remarkably consistent across different concepts, as evidenced by the famous analogy:
$\bm{v}(\text{king})-\bm{v}(\text{man})+\bm{v}(\text{woman})\approx\bm{v}(\text{queen}),$ (1)
This suggests that abstract features, like “gender”, correspond to specific directions within the high-dimensional embedding space.
A critical question thus arises: how many distinct semantic features can be encoded within an embedding space of dimension $d$? If we demand that each feature corresponds to a perfectly unique, and therefore mathematically orthogonal, direction, the capacity of the space is strictly limited. By definition, a $d$-dimensional space can contain at most $d$ mutually orthogonal vectors. For any two such vectors $\bm{v}_{i}$ and $\bm{v}_{j}$, their dot product must be zero ($\bm{v}_{i}\cdot\bm{v}_{j}=0$). This linear scaling would severely constrain the model’s expressive power.
However, suppose the requirement for perfect orthogonality can be relaxed. Instead of perfect distinction, we can consider features to be sufficiently distinct if their corresponding vectors are “almost” orthogonal, meaning their dot product is very close to zero ($|\bm{v}_{i}\cdot\bm{v}_{j}|\leq\varepsilon$ for some small $\varepsilon>0$). This relaxation from strict orthogonality to quasi-orthogonality leads to a dramatic, non-intuitive increase in the expressive capacity of the space.
The theoretical underpinning for this phenomenon is the Johnson-Lindenstrauss lemma [7]. The lemma states that a set of $n$ points in a high-dimensional space $\mathbb{R}^{d}$ can be projected into a much lower-dimensional space $\mathbb{R}^{k}$, where $k=O(\varepsilon^{-2}\log n)$, while preserving all pairwise distances to within a factor of $(1\pm\varepsilon)$. Indeed, this bound on the target dimension of this embedding is tight [8, 9].
This preservation of distances translates into the preservation of inner products of the respective position vectors. And the (almost) vanishing of inner products indicates (almost) orthogonality. By applying the (real) polarization identity, $\bm{u}\cdot\bm{v}=\frac{1}{2}\left(\|\bm{u}\|^{2}+\|\bm{v}\|^{2}-\|\bm{u}-\bm{v}\|^{2}\right)$, the preservation of the Euclidean norms of the position vectors and their differences—and thus the distances of the respective points—ensures that pairwise inner products are also preserved to within a small additive error. This preservation of inner products is the crucial property for establishing quasi-orthogonality.
While often cited for dimensionality reduction, the tight Johnson-Lindenstrauss lemma’s implications for the capacity of a high-dimensional space are profound: Inverting this finding yields an exponential packing phenomenon. Let $N_{\varepsilon}(d)$ be the largest size of a set of unit vectors in $\mathbb{R}^{d}$ with pairwise inner products bounded by $\varepsilon$ in magnitude. Then there exist universal constants $c,C>0$ such that
$\exp\!\bigl(c\,\varepsilon^{2}d\bigr)\;\leq\;N_{\varepsilon}(d)\;\leq\;\exp\!\bigl(C\,\varepsilon^{2}d\bigr).$
That is, pointedly stated, a $d$-dimensional space can accommodate a set of approximately $e^{O(d\varepsilon^{2})}=e^{\varepsilon^{2}O(d)}$ such vectors. This exponential growth provides a mathematical basis for the remarkable ability of LLMs to encode a vast and nuanced landscape of semantic features, far exceeding the linear bounds of the space’s dimensionality: If each interpretable property corresponds to a direction and decoders tolerate pairwise correlations up to $\varepsilon$, then a $d$-dimensional space can in principle host on the order of $e^{\varepsilon^{2}O(d)}$ property directions with bounded interference. We refer to this exponentially larger supply of usable, nearly independent directions as quasi-dimensionality.
VIII Polysemy
@@
VIII.1 Dynamic Contextualization via Self-Attention
@@
VIII.2 An Alternative Framework: Quantum Contextuality
@@
IX The Origin of Stochasticity in responses
@@
X Algorithmic Capacity and the Role of Stochastic Seeding
@@ X.1 The LLM as a Formal System
X.2 Stochasticity as an Engine for Discovery
@@
X.3 Comparison to Genetic Algorithms
@@
XI Comparison to Quantum Mechanics
@@
XI.1 Vectors, Inner Products, and Contexts
@@
XI.2 Dynamics: Unitary Evolution vs. Transformer Block Updates
@@
XI.3 Contextuality in Both Senses
@@
XI.4 Operationalizing an Internal Observer
@@
XI.5 Orthogonal Bases and Functional Decomposition
@@
XI.6 Analogies, Disanalogies, and Conclusion
@@ XII Quantum Attention
XII.1 Goals of Quantum Attention
@@
XII.2 Constraints and Learning Mechanisms
@@
XII.3 Comparison with Classical Attention
@@ Appendix C The Self-Attention Mechanism
Let us consider a sequence of $n$ input tokens—say, an “unfinished” sentence—to be finished by the token derived from the output $Y$ in Eq. (43). Each token is represented by a learned embedding vector, and these row vectors are stacked to form an input matrix $X\in\mathbb{R}^{n\times d}$, where $d$ is the dimension of the embedding space (often called $d_{\text{model}}$ in literature). The self-attention mechanism computes an output matrix $Y\in\mathbb{R}^{n\times d}$ where each output row $\bm{y}_{i}$ is a new, contextualized representation of the corresponding input token $\bm{x}_{i}$. This is achieved through the following steps.
@@ C.1 Projection to Query, Key, and Value
First, the input matrix $X$ is linearly projected into three distinct matrices: Query ($\bm{Q}$), Key ($\bm{K}$), and Value ($\bm{V}$). This is accomplished using three learned weight matrices: $\bm{W}_{Q}\in\mathbb{R}^{d\times d_{k}}$, $\bm{W}_{K}\in\mathbb{R}^{d\times d_{k}}$, and $\bm{W}_{V}\in\mathbb{R}^{d\times d_{v}}$. Typically, for the single-head attention block, the dimensions are set such that $d=d_{k}=d_{v}$ [2] (for multi-head attention with $h$ parallel heads discussed in D.1, $d_{k}=d_{v}=d/h$).
@@ C.4 Output Computation
A loss function is calculated to quantify the divergence between the predicted token distribution and the ground-truth token—the target that is the actual token completion—from the training set. The gradient of this loss is then backpropagated through the model to adjust all learnable parameters, including the embedding matrix and the query ($W_{Q}$), key ($\bm{W}_{K}$), and value ($\bm{W}_{V}$) matrices. (For multi-head attention below, this includes the output matrix $\bm{W}_{O}$.)
@@ D.1 Multi-Head Attention
    Projection: Instead of one set of projection matrices, the model learns $h$ independent sets of matrices ($\bm{W}_{Q,i},\bm{W}_{K,i},\bm{W}_{V,i}$ for each head $i=1,\dots,h$), with $d_{k}=d_{v}=d/h$. The input $X$ is projected through each of these sets, creating $h$ different Query, Key, and Value representations.
@@
    Concatenation and Final Projection: The outputs of these $h$ heads are “horizontally” concatenated along the feature dimension into a single, larger matrix. This matrix is then projected back to the original model dimension $d$ using an additional learned weight matrix, $\bm{W}_{O}\in\mathbb{R}^{hd_{v}\times d}$, which learns to aggregate the information from all heads.
    $\text{MultiHead}(X)=\text{Concat}[\text{head}_{1},\dots,\text{head}_{h}]\bm{W}_{O}.$ (46)